OpenStudy (anonymous):

1,567,000 = 925,000 × (1 + x)^2 , I don't know what I'm doing wrong , I keep getting .83 but the answer is around .3

myininaya (myininaya):

solving for x?

myininaya (myininaya):

those numbers are big

OpenStudy (lilg132):

i think it may be 1,567,000/925,000 = (1 +x)^2

OpenStudy (anonymous):

divide both sides by 925000, then take the square root of both sides, then subtract 1 and you get .30

OpenStudy (anonymous):

Yes , step by step would be nice :)

OpenStudy (lilg132):

(1,567,000/925,000) - 1^2 = x^2

myininaya (myininaya):

\[\frac{1567000}{925000}=(1+x)^2\] \[\frac{1567}{925}=(1+x)^2\]

myininaya (myininaya):

now take square root of both sides

myininaya (myininaya):

\[\sqrt{\frac{1567}{925}}=\sqrt{(1+x)^2}\] recall that \[\sqrt{(1+x)^2}=|1+x|=\pm (1+x)\] so we have \[\pm \sqrt{\frac{1567}{925}} =1+x\]

myininaya (myininaya):

now subtract 1 on both sides \[-1 \pm \sqrt{\frac{1567}{925}}=x\]

myininaya (myininaya):

\[x=-1 \pm \sqrt{\frac{1567}{925}}\]

OpenStudy (anonymous):

1567/925= 1.694, so sqrt(1.694)=sqrt((1+x)^2) so sqrt(1.694)=(1+x) so sqrt(1.694) - 1 = x 1.30-1=x x=.3

myininaya (myininaya):

thats one answer

myininaya (myininaya):

-2.3 is the other

myininaya (myininaya):

.3,-2.3 are approximations just so you know i gave you the exact values above